OpenStudy (swaqqout_kid1):

What is the slope of the line represented by the equation ? (Points : 1) –5 6

OpenStudy (swaqqout_kid1):

A-5 B 6 C 5/6 D 1/6

OpenStudy (swaqqout_kid1):

Q:Y=x/6-5

OpenStudy (anonymous):

The standard equation: y=mx+b where m coefficient of slope and b is the free number (y intercept) They already gave you the standard form, y=x/6-5

OpenStudy (anonymous):

c

OpenStudy (anonymous):

So for example, for this equation: y=2x-1 The slope would be 2.

OpenStudy (anonymous):

(It's not C by the way)

OpenStudy (anonymous):

yes it is u stupid hoe

OpenStudy (swaqqout_kid1):

so its C

OpenStudy (anonymous):

What's the coefficient of x?

OpenStudy (swaqqout_kid1):

idk

OpenStudy (anonymous):

Example: x/2 = 1/2 times x

OpenStudy (anonymous):

What about x/6?

OpenStudy (anonymous):

\[\frac{ x }{ 6 } = x\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):



OpenStudy (swaqqout_kid1):

its D

OpenStudy (anonymous):

Yup.